# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v2

@@ Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the connections among numerous existing methods to gain in-depth understanding of this field. To bridge these gaps, in this paper, we timely conduct an extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
@@
To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution flatter and more suitable for quantization. Quantization error mitigation, on the other hand, involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry. Third, we also timely analyze and evaluate the latest MXFP4 and NVFP4 data format and its performance.
@@
Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Second, as the granularity becomes finer, the performance improves, but this also incurs additional storage overhead. We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations. Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on the performance. We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
keywords:
Model Quantization, Post-training Quantization , LLM
1 Introduction
@@
Large language models (LLMs) have achieved remarkable performance on diverse tasks and are increasingly being applied in various scenarios. Nonetheless, LLMs inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on source-limited GPUs.
Post-training quantization (PTQ) is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. LLMs quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [1, 2, 3, 4], W8A8 [5, 6, 7] and W4A16 [8, 9, 10, 11]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively. Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed. However, applying lossless W4A4 quantization to LLMs remains a challenging problem. Thus, we concentrate our efforts on exploring W4A4 quantization.
@@
In this paper, we mainly investigate three dimensions. First, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. OSTQuant [8] and FlatQuant [12] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [8], a gradient-based method for mitigating quantization error, to quantize the transformed weights. Besides, many techniques share similar transformations, for instance, Quip# [10] and QuaRot [13] both use rotation, also, AWQ [11] and SmoothQuant [6] both use scaling. To achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
@@
Third, NVIDIA’s GeForce RTX 50 Series, which is newly released and based on the Blackwell architecture, includes support for the FP4 data format (e.g., NVFP4 and MXFP4) in its Tensor Cores, showing significant advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization using the popular methods in INT4, showing the great timeliness of our research. For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on the performance.
2 Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section 2.1 in terms of two steps: pre-quantization transformation and quantization error mitigation. Table 1 shows our decomposition of the popular quantization methods. Section 2.2 focuses on settings, including symmetry and granularity of quantization methods. Meanwhile, Table 2 summarizes the settings of existing methods, showing that they are not compared on the same ground. Section 2.3 discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
@@ 2 Review on Existing Quantization Methods
GPTQ [8] - GPTQ
ZeroQuant-v2 [14] - Low-rank
@@
Table 1: Two steps decomposition of quantization methods.
2.1 Two-steps of quantization
2.1.1 Pre-quantization transformation
Methods Precision Symmetric Granularity
@@ 2.1.1 Pre-quantization transformation
GPTQ [8] W3A16,W4A16 T - G
ZeroQuant-v2 [14] W4A16,W3A16,W2A16,W4A8 T,F T,F G
QuIP [9] W2A16,W3A16,W4A16 T - G
AWQ [11] W3A16,W4A16 T - G
@@
Table 2: Symmetry and granularity of quantization methods. In granularity, T stands for per-tensor, C represents per-channel or per-token, and G means per-group.
@@
Table 3: Comparison of shifting-based quantization methods.
As shown in Table 3, Outlier Suppression+ [7] uses calibration to get the vector $T$, while OmniQuant [2] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
@@
Methods Scaling Strategy Scaling parameters $S$
SmoothQuant [6] Scaling only $\begin{aligned} &S_{i}=\frac{\max(|X_{i}|)^{\alpha}}{\max(|W_{i}|)^{1-\alpha}}\end{aligned}$
AWQ [11] Scaling only $\begin{aligned} &S_{i}=|X_{i}|_{avg}^{\alpha}\\ &\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})\end{aligned}$
OS+ [7] Post-shifting scaling $\begin{aligned} &S_{i}=\max\left(1.0,\frac{\max(X_{i}-T_{i})}{t}\right)\\ &t^{*}=\arg\min_{t}\mathcal{L}(S_{i})\end{aligned}$
OmniQuant [2] Post-shifting scaling Gradient-based optimization
FlatQuant [12] Pre-rotation scaling Calibration initialization and gradient-based optimization
OSTQuant [4] Post-rotation scaling Gradient-based optimization
Table 4: Comparison of scaling-based quantization methods.
As shown in Table 4, SmoothQuant [6] uses calibration to compute $S$. AWQ [11], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights. Its scaling parameter is determined by the average magnitude of the activation, while the optimal exponent $\alpha^{*}$ is found by minimizing the loss through grid searching. Outlier Suppression+ [7] also uses grid searching optimization to find the optimal scaling after applying shifting. Unlike adopting the hand-crafting parameters in calibration, Outlier Suppression+ [7] uses grid searching optimization to find the optimal scaling after applying shifting. OmniQuant [2] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error. FlatQuant [12] and OSTQuant [4] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings.
@@
Methods Rotation Strategy Rotation parameters $O$
@@
FlatQuant [12] Post-scaling Rotation Gradient-based optimization
OSTQuant [4] Pre-scaling Rotation Gradient-based optimization
Table 5: Comparison of rotation-based quantization methods.
As shown in Table 5, QuIP [9] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [10] replaces this with random Hadamard matrices for greater efficiency. QuaRot [13] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [3] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
@@
2.1.2 Quantization error mitigation
@@
$\Delta\mathcal{L}=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)\approx\frac{1}{2}(dq(W_{q})-W)^{\top}H(dq(W_{q})-W)$ (4)
where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$. $W_{q}$ and $dq(W_{q})$ denote the quantized weights and their dequantized version, respectively. Self-compensation methods need to find an optimal weight perturbation $\delta$ to add to $W$, such that the quantization loss error in Eq. (4) is minimized. Table 6 summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
@@
Table 6: Comparison of self-compensation quantization methods.
@@
Low-rank decomposition compensation
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to error correction.
@@
where $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error. Table 7 illustrates the computation process of the low-rank terms in some quantization methods.
@@
ZeroQuant-v2 [14] $\text{SVD}(dq(W_{q})-W)$
LQER [16] $\text{SVD}(dq(W_{q})-W)$
L2QER [16] $\text{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
@@
QERA [22] $\text{SVD}(R^{\frac{1}{2}}_{X\!X}(dq(W_{q})-W)),A:=(R^{\frac{1}{2}}_{X\!X})^{-1}A$
Table 7: Comparison of low-rank decomposition compensation quantization methods. $A,B$ represents the low rank terms. $R_{X\!X}$ denotes the autocorrelation matrix of the input $X$.
ZeroQuant-v2 [14] and LQER [16] apply SVD composition to reconstruct the quantization error matrix $E=W-W_{q}$. Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$. This operation aims to more precisely approximate the quantization error corresponding to the salient weights.
@@
2.2 Settings of quantization
2.2.1 Symmetric vs. asymmetric
Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values.
(a) Symmetric quantization
(b) Asymmetric quantization
@@ 2.2 Settings of quantization
Symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy. As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
2.2.2 Granularity of quantization

### v3

@@ Abstract
For large language models (LLMs), post-training quantization (PTQ) can significantly reduce memory footprint and computational overhead. Model quantization is a rapidly evolving research field. Though many papers have reported breakthrough performance, they may not conduct experiments on the same ground since one quantization method solution usually contains multiple components. In addition, it is also important to analyze the connections among numerous existing methods to gain an in-depth understanding of this field. To bridge these gaps, in this paper, we conduct an up-to-date and extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. To our knowledge, this fair and extensive investigation remains critically important yet significantly underexplored in the research community.
@@
To better understand the connections, first, we decouple the published quantization methods into two steps: pre-quantization transformation and quantization error mitigation. We define the former as a preprocessing step applied before quantization to reduce the impact of outliers, making the data distribution more suitable for quantization. Quantization error mitigation involves techniques that offset the errors introduced during quantization, thereby enhancing model performance. Second, we extensively evaluate and analyze the impact of different settings of quantization methods, including granularity and symmetry. Third, we also analyze and evaluate the latest MXFP4 and NVFP4 data formats and their performance.
@@
Our experimental results first demonstrate that the optimized rotation and scaling yield the best performance for pre-quantization transformation, and combining low-rank compensation with GPTQ occasionally outperforms using GPTQ alone for quantization error mitigation. Second, as the granularity becomes finer, the performance improves, but this also incurs additional storage overhead. We also verified that applying asymmetric quantization to weights yields much smaller benefits compared to applying it to activations, which is consistent with the common practice of using asymmetric quantization primarily to activations. Third, we explore the potential of the latest FP4 quantization and find that the format and precision of scaling factors have a significant impact on performance. We also show the performance of the optimal pre-quantization transformation strategy for INT4 on FP4, revealing that rotation-based methods demonstrate little improvement in performance for MXFP4 and NVFP4, inspiring further investigation in this field.
I Introduction
@@
Large language models (LLMs) have achieved remarkable performance on diverse tasks and are increasingly being applied in various scenarios. Nonetheless, LLMs’ inference process incurs significant memory and computational overhead, leading to high inference latency and posing challenges for computations on resource-limited GPUs.
Quantization is a promising solution to reduce the memory and computational cost of LLMs, enhancing their inference efficiency. In recent years, the quantization techniques have developed rapidly. Various approaches have been proposed, including Quantization-Aware Training (QAT), Post-Training Quantization (PTQ), and hybrid methods that combine quantization with knowledge distillation or pruning. In this work, we focus solely on quantization, without considering hybrid approaches. In addition, we focus on PTQ, because it is more widely used than QAT in real applications.
Additionally, LLMs’ quantization has been studied in a wide spectrum of precision configurations. Representative settings include W4A4 [43, 32, 22, 13], W8A8 [8, 38, 37] and W4A16 [10, 6, 35, 19] and mixed-precision settings such as W4A8 [20]: W4A4 quantizes both weights and activations to 4-bit precision (INT4); W8A8 and W4A16 quantize weights to INT8 and INT4 and activations to INT8 and FP16/BF16 respectively; Mixed-precision methods aim to improve accuracy by using higher bit-widths for parts that are more difficult to quantize [43].
Among these three settings, W4A4 is currently the main focus of research, as it holds great potential for improving memory efficiency and inference speed. However, applying lossless W4A4 quantization to LLMs remains a challenging problem. To make our evaluation more cutting-edge and provide the research community with up-to-date insights, we select W4A4 as the primary precision configuration for our study.
Many papers report breakthrough performance in quantization; however, their findings are often not directly comparable, as quantization often contains multiple configurations. It is important to evaluate these methods on the same ground.
@@
Besides, we find that existing methods are often combinations of multiple techniques, since using a single technique alone typically fails to achieve excellent quantization performance. OSTQuant [10] and FlatQuant [34] employ multiple techniques to adjust the weights and activations, including rotation and scaling, and also apply the GPTQ [10], a gradient-based method for mitigating quantization error, to quantize the transformed weights. Besides, many techniques share similar transformations, for instance, Quip# [35] and QuaRot [3] both use rotation, also, AWQ [19] and SmoothQuant [38] both use scaling. Analysis of the interconnections among diverse methods is critical for a thorough understanding of the field.
To bridge these gaps, in this paper, we conduct an up-to-date and extensive review of the state-of-the-art methods and also perform comprehensive evaluations on the same ground to ensure fair comparisons of these methods. Although several excellent survey papers on quantization exist [17, 14], this work makes more extensive evaluations and categorizes more state-of-the-art quantization methods from different perspectives.
We mainly investigate three dimensions. First, to achieve a better understanding of existing methods and make our evaluation on the same ground, we decompose the quantization process into two key steps: pre-quantization transformation and quantization error mitigation. The pre-quantization transformation step refers to preprocessing the data before quantization in order to mitigate the effect of outliers, making the data distribution smoother and more quantization-friendly. The quantization error mitigation step involves applying techniques to compensate for the errors introduced by quantization, thereby reducing the overall quantization error and improving performance. For each step, we select several state-of-the-art techniques and conduct an evaluation of the effectiveness of their combinations.
@@
Third, NVIDIA’s GeForce RTX 50 Series, which is newly released and based on the Blackwell architecture, includes support for the FP4 data format (e.g., NVFP4 and MXFP4) in its Tensor Cores, showing significant advantages over INT4. However, the research on FP4 has not been well explored. To bridge this gap, we conduct research on weight-activation FP4 quantization under our two-step quantization framework in INT4, showing the great timeliness of our research. For FP4 quantization, we also observe that the format and precision of scaling factors have a significant impact on performance.
II Review on Existing Quantization Methods
In this section, we first review recent quantization algorithms in Section II-A in terms of two steps: pre-quantization transformation and quantization error mitigation. Table I shows our decomposition of the popular quantization methods. Section II-B focuses on settings, including symmetry and granularity of quantization methods. Meanwhile, Table II summarizes the settings of existing methods, showing that they are not compared on the same ground. Section II-C discusses FP4 quantization and its variants, including MXFP4 and NVFP4.
TABLE I: Two-step decomposition of quantization methods.
@@ II Review on Existing Quantization Methods
GPTQ [10] GPTQ
ZeroQuant-v2 [39] Low-rank
@@
TABLE II: Symmetry and granularity of quantization methods. For granularity, T: per-tensor, C: per-channel/token, G: per-group.
Methods Precision Symmetric Gran.
@@
GPTQ [10] W3A16, W4A16 T G
ZeroQuant-v2 [39] W2/3/4A16, W4A8 T, F T, F G
QuIP [6] W2/3/4A16 T G
AWQ [19] W3A16, W4A16 T G
@@
II-A Two steps of quantization
II-A1 Pre-quantization transformation
@@
TABLE III: Comparison of shifting-based quantization methods.
@@
As shown in Table III, Outlier Suppression+ [37] uses calibration to get the vector $T$, while OmniQuant [32] employs learnable channel-wise shifting for outlier suppression, combined with the learnable scaling technique discussed in the next paragraph, and jointly optimizes these two kinds of pre-quantization transformation parameters.
@@
TABLE IV: Comparison of scaling-based quantization methods.
Methods Strategy Scaling parameters $S$
SmoothQuant [38] Scaling only $S_{i}=\max(|X_{i}|)^{\alpha}/\max(|W_{i}|)^{1-\alpha}$
AWQ [19] Scaling only $S_{i}=|X_{i}|_{\text{avg}}^{\alpha}$ $\alpha^{*}=\arg\min_{\alpha}\mathcal{L}(S_{Xi}^{\alpha})$
OS+ [37] Post-shift $S_{i}=\max(1.0,\frac{\max(X_{i}-T_{i})}{t})$ $t^{*}=\arg\min_{t}\mathcal{L}(S_{i})$
OmniQuant [32] Post-shift Gradient-based optimization
FlatQuant [34] Pre-rot. Calibration init. & Grad-based
OSTQuant [13] Post-rot. Gradient-based optimization
As shown in Table IV, SmoothQuant [38] uses calibration to compute $S$. AWQ [19], as a weight-only quantization, approaches the problem from the perspective of protecting salient weights. Its scaling parameter is determined by the average magnitude of the activation, while the optimal exponent $\alpha^{*}$ is found by minimizing the loss through grid searching. Outlier Suppression+ [37] also uses grid searching optimization to find the optimal scaling after applying shifting. OmniQuant [32] optimizes the quantization parameters via gradient-based methods to minimize the block-wise quantization error. FlatQuant [34] and OSTQuant [13] both use learnable per-channel scaling, but at different stages of pre-quantization transformation, leading to different practical meanings.
@@
TABLE V: Comparison of rotation-based quantization methods.
Methods Strategy Rotation parameters $O$
@@
FlatQuant [34] Post-scaling Gradient-based optimization
OSTQuant [13] Pre-scaling Gradient-based optimization
As shown in Table V, QuIP [6] first introduced rotation with random orthogonal matrices to reduce incoherence in weight-only quantization. QuIP# [35] replaces this with random Hadamard matrices for greater efficiency. QuaRot [3] extends rotation to both weights and activations quantization using random Hadamard matrices, effectively eliminating outliers. SpinQuant [22] further improves stability by optimizing the rotation matrix on the Stiefel manifold to minimize quantization-induced performance variance, while still defaulting to random Hadamard for a few online rotations.
@@
II-A2 Quantization error mitigation
@@
$\Delta\mathcal{L}$ $=\mathcal{L}(dq(W_{q}))-\mathcal{L}(W)$ (4)
$\approx\frac{1}{2}\bigl(dq(W_{q})-W\bigr)^{\top}H\bigl(dq(W_{q})-W\bigr)$
where $\mathcal{L}(W)$ is the loss function of the model evaluated at weights $W$. $W_{q}$ and $dq(W_{q})$ denote the quantized weights and their dequantized version, respectively. Self-compensation methods need to find an optimal weight perturbation $\delta$ to add to $W$, such that the quantization loss error in Eq. (4) is minimized. Table VI summarizes the computation strategies of $\delta$ adopted by various self-compensation quantization methods.
TABLE VI: Comparison of self-compensation quantization methods.
@@
Low-rank compensation
Beyond compensating errors directly on the original weights, an alternative approach is to introduce an auxiliary branch dedicated to quantization error mitigation.
@@
where $A\in\mathbb{R}^{C_{in}\times r}$, $B\in\mathbb{R}^{r\times C_{out}}$ are low-rank matrices, and $r\ll\min(d,k)$ is the bottleneck rank. The product $AB\in\mathbb{R}^{d\times k}$ serves as a low-rank approximation of the quantization error. Table VII illustrates the computation process of the low-rank terms in some quantization methods.
TABLE VII: Comparison of low-rank compensation quantization methods.A,BA,B represent the low-rank terms. $R_{XX}$ denotes the autocorrelation matrix of input $X$.
@@
ZeroQuant-v2 [39] $\mathrm{SVD}(dq(W_{q})-W)$
LQER [41] $\mathrm{SVD}(dq(W_{q})-W)$
L2QER [41] $\mathrm{SVD}(S(dq(W_{q})-W)),A:=S^{-1}A$
@@
QERA [42] $\mathrm{SVD}(R_{XX}^{1/2}(dq(W_{q})-W)),$ $A:=(R_{XX}^{1/2})^{-1}A$
ZeroQuant-v2 [39] and LQER [41] apply SVD to reconstruct the quantization error matrix $E=W-dq(W_{q})$. Based on LQER, L2QER, which is proposed in LQER, scales the $E$ before applying SVD heuristically, and inversely applies scales in $A_{k}$. This operation aims to more precisely approximate the quantization error corresponding to the salient weights.
@@
II-B Settings of quantization
II-B1 Symmetric vs. asymmetric
Uniform quantizers can be broadly categorized into symmetric and asymmetric schemes, depending on how the zero point is handled during the mapping from floating-point to integer values, as illustrated in Fig. 3.
((a))
((b))
@@ II-B Settings of quantization
Symmetric quantization enables a very simple dequantization process. For linear layers, dequantization can be efficiently performed by multiplying the quantized output by the product of the scale factors for the weights and activations. This simplifies computation and is often preferred in hardware implementations. However, symmetric quantization may lead to suboptimal representation when the data distribution is not centered around zero.
Asymmetric quantization incurs additional computational cost during dequantization, while it often outperforms symmetric quantization in terms of accuracy.
As such, the selection of quantization symmetry should be made with consideration of the specific use case and performance-complexity trade-offs.
II-B2 Granularity of quantization